# Improving Visual Place Recognition Based Robot Navigation Through Verification of Localization Estimates

### v1

@@ Abstract
Visual Place Recognition (VPR) systems often have imperfect performance, which affects robot navigation decisions. This research introduces a novel Multi-Layer Perceptron (MLP) integrity monitor for VPR which demonstrates improved performance and generalizability over the previous state-of-the-art SVM approach, removing per-environment training and reducing manual tuning requirements. We test our proposed system in extensive real-world experiments, where we also present two real-time integrity-based VPR verification methods: an instantaneous rejection method for a robot navigating to a goal zone (Experiment 1); and a historical method that takes a best, verified, match from its recent trajectory and uses an odometer to extrapolate forwards to a current position estimate (Experiment 2). Noteworthy results for Experiment 1 include a decrease in aggregate mean along-track goal error from $\approx\!9.8m$ to $\approx\!3.1m$ in missions the robot pursued to completion, and an increase in the aggregate rate of successful mission completion from $\approx\!41\%$ to $\approx\!55\%$. Experiment 2 showed a decrease in aggregate mean along-track localization error from $\approx\!2.0m$ to $\approx\!0.5m$, and an increase in the aggregate precision of localization attempts from $\approx\!97\%$ to $\approx\!99\%$. Overall, our results demonstrate the practical usefulness of a VPR integrity monitor in real-world robotics to improve VPR localization and consequent navigation performance.
@@ I Introduction
Figure 1: Overview of our system, demonstrating how our addition of a predictive verification system (specifically, we train a Multi-Layer Perceptron (MLP) network) results in safer navigation. See Figure 2 for a more detailed diagram of the proposed integrity-based verification system.
@@
    We present a novel approach using a Multi-Layer Perceptron network for VPR integrity monitoring, rather than the previously adopted Support Vector Machines, improving performance, generalizability across datasets, and outcomes in robot navigation paradigms.
@@
The letter proceeds as follows: in Section II we review the literature, in Section III we detail our approach, in Section IV we introduce our experimental methods, and in Section V we discuss the results.
@@ II-A Visual Place Recognition
Typically VPR involves extracting features from a current (query) image and comparing these with features from a set of reference images taken along a previously-traversed path to determine a best-match[3]. Numerous VPR techniques exist, with feature extraction and matching varying in complexity and performance [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15], noting the recent techniques MixVPR [16], DinoV2 SALAD [17], and AnyLoc [18].
When deployed on real-world navigating robots, VPR is susceptible to potentially large localization errors due to varied environmental conditions, lighting, and presence of dynamic objects [19]. Although stand-alone systems have no match validation, some research has attempted to characterize the uncertainty. Uncertainty is often modelled using known distributions, however, VPR errors are discontinuous and non-Gaussian [20]. Sequence matching [21, 22, 23], Bayesian techniques [24, 25, 26, 27, 28] and particle filters [28, 29, 30] may be used to help reduce error but adds cost in complexity and sensing. Some research has focused on predicting which VPR combinations will perform best in a given environment [31, 32], but these methods are computationally intensive.
@@ II-B Localization Integrity
Some integrity monitoring approaches have been proposed for stereo visual odometry [34, 35, 36, 37]; or for fusing visual odometry with sensors such as GNSS, LiDAR and radar [38, 39, 40, 41]. Other approaches include vision-based lane detectors [42], landmarks localization [43], or objects of interest localization within a scene [39]. However, none of these methods specifically address VPR. In our previous work [44], we proposed SVM-based integrity prediction methods for VPR such that each match’s accuracy is predicted to be within a tolerance; but this was investigated passively on datasets, not actively navigating robots. Regardless of method, integrating an integrity monitor is an attractive proposition as it can enable a robot to introspectively improve real-time operational decisions.
@@ III Approach
In this section we describe the proposed instantaneous and historical localization methods with their requirements.
@@ III-A VPR Technique Requirements
The integrity monitor we build on from [44] only requires some form of distance vector input for predictions, making it compatible with many, if not all, VPR techniques. The distance vectors consist of mathematical distances between query image features and those from a reference set; we refer to these distances as ‘match distances’. In this work, we maintain compatibility with all, or at least the vast majority of, VPR techniques by similarly using information already available from the VPR process to verify VPR integrity. We demonstrate this by testing using various VPR techniques, including AP-GeM [45], NetVLAD [5], and SALAD [17].
@@ III-B Integrity Prediction
We improve on the previously adopted SVM model from [44] by implementing a small multi-layer perceptron (MLP) network to utilise the increased learning capability of neural networks whilst maintaining the light-weight nature of the predictor. We use a total of 192 hand-crafted statistical features from four vectors, 48 per vector, as inputs for classification. These vectors include the distance vector as per [44], as well as the query features, the best-match reference features, and the difference between these query and reference features. For brevity we include a simplified visualization of the process (Fig. 2), but the full mathematical formulation of these features can be found in the mentioned repository1.
Figure 2: An overview of how inputs for the MLP integrity monitor are extracted. For a given query, we take the match distance vector, query feature vector, top reference match feature vector, and the difference of these VPR feature vectors. Each vector goes through a statistical feature extractor, and the results are concatenated before passing into the integrity monitor.
@@ III-C MLP Integrity Monitor Training
For a given query image with index $k$, the input vector for the MLP network becomes a 192-dimensional vector created through concatenating four groups of 48 statistical features. These groups are calculated from the $k^{\text{th}}$: 1) distance vector, $\mathbf{D}$ [^7] Bold denotes arrays. Subscripts denote indices, which start from $1$., 2) query VPR feature vector, $\mathbf{Q}$, 3) top VPR match reference feature vector, $\mathbf{R}$, and 4) VPR feature difference vector, $\mathbf{V}=\mathbf{R}-\mathbf{Q}$. The query training label, $\mathbf{P}_{k}$, simply becomes a binary value indicating whether the top VPR match corresponds to a reference in-tolerance, $1$, or out-of-tolerance, 00. We formulate the problem as a regression task using a weighted mean-squared-error (MSE) loss, providing control over the rate queries are predicted as out-of-tolerance. The loss function, $L(\mathbf{P},\mathbf{\hat{P}},\alpha)$, is defined as:
@@
where $\mathbf{\hat{P}}_{k}$ is the integrity prediction from the MLP network, $\alpha$ is a scalar weight, and $N$ is the dimension of $\mathbf{P}$ and $\mathbf{\hat{P}}$. We set $\alpha>1$ to produce a cautious integrity monitor that reduces the number of out-of-tolerance VPR matches that are incorrectly predicted as in-tolerance (false positives), given these pose a greater risk to robot navigation tasks.
III-D Localizing with VPR and Integrity
In the instantaneous scenario where each query is treated independently, the VPR system performs query-to-reference matching and provides a position estimate using the retrieved 2D pose ($x$,$y$,$\theta$) from where the best-match reference image was collected. The integrity monitor (Section III-B) then verifies the unprocessed output by only accepting position estimates that are predicted to be in-tolerance of the true position; all predicted out-of-tolerance estimates are rejected.
III-E Utilizing Historical Data
We can improve the instantaneous localization decisions (Section III-D) by utilizing an odometer to extrapolate a position estimate based on a recent history of predicted in-tolerance VPR matches. The use of historical information is a well-known component of many navigation systems: our key contribution here is the addition of a precursor step where all entries in the history are verified (such that only predicted in-tolerance matches are retained), distinguishing our method from those which only perform statistical or threshold filtering. If the system finds at least one verified historical match, then it will proceed to localize using that match and recent odometer measurements. We diagram this process in Figure 3.
Figure 3: Historical method: we rank a recent history of VPR matches by their match distance (1), reduce to verified matches (2), then extrapolate from the best, verified match to estimated position using the odometer difference (3).
This process can be described by the following. Consider some unknown current vehicle pose $\chi$ and known odometer scalars $\omega$. Within the previous $d$ meters, we sample VPR match indices $\mathbf{M}$ and odometer scalars $\mathbf{W}$, collected together at irregular spatiotemporal steps. For each match index $\mathbf{M}_{i}$, we find a corresponding: match distance $\mathbf{D}_{i}$, integrity prediction $\mathbf{\hat{P}}_{i}$, and pose estimate $\mathbf{X}_{\mathbf{M}_{i}}\in\mathbf{X}$, where $\mathbf{X}$ is the ordered set of 2D poses with elements indexed to each reference image (dimension $n$). We assume $\chi\equiv\mathbf{X}_{q}$, where $q$ is the index of the closest (Euclidean) pose in $\mathbf{X}$ to $\chi$. Using these as the recent history contents ($\mathbf{M},\mathbf{D},\mathbf{W},\mathbf{\hat{P}}$, and $\mathbf{X}_{\mathbf{M}}$, per Figure 3), we now work to determine $q$.
@@ III-E Utilizing Historical Data
We then compute an array of the running sum of the elements of $\mathbf{T}$, denoted as $\mathbf{S}$, from which we find $q$ and thus $\chi$ per Figure 3, Step 3:
@@ IV Experimental Procedure
In this section, we introduce our two experiment designs for assessing the performance of our proposed instantaneous and historical methods, as well as our parameters, platform, data collection scheme, and techniques for comparison. We begin this section importantly by stating that in the following described experiments we were able to train a single MLP integrity monitor for each VPR technique to perform across all experiments, test datasets and conditions.
@@ IV-A Experiment 1 Design
In the first experiment scenario, a robot moves along a path whilst attempting to autonomously determine if it has reached an end-goal zone. We implement the method introduced in Section III-D for localization. The mission is successfully completed if the robot correctly identifies it has arrived at the end-goal within an assessment tolerance of $\pm$0.5m, otherwise the mission is failed, as shown in Figure 4.
@@
For fair assessment, given the underlying baseline system performance changes along the route due to environmental and dynamic variations, we randomly select 50 start locations over the reference path, with end-goal locations allocated at 5m, 10m, 25m and 50m from each starting position. As navigation decisions require a tighter tolerance, the robot will flag arrival if it estimates its location is anywhere beyond the point 10cm before the end-goal center. This provides a margin of error in achieving the looser, actual, target tolerance.
@@ IV-B Experiment 2 Design
In the second experiment scenario, a robot repeats a pre-mapped traverse whilst localizing using the method introduced in Section III-E. In this scenario, the goal is to maximize precision. We select a history length of 1.5m. For fair analysis the robot must initialize the history by traversing this distance.
@@ IV-D Data Collection: Environments and Conditions
We conducted both experiments in two different environments: Office, containing indoor images from the QUT Centre for Robotics; and Campus, containing outdoor images from the QUT Gardens Point campus. For both environments we recorded two sets of conditions: Normal, where there is a typical presence of adversity given the environment (some minor lighting variation and dynamic or static objects); and Adverse, where there is a heightened presence of adversity, such as increased lighting variation (morning to evening), and more dynamic or static objects. Data capture periods were selected to control these introduced adversities. We emphasize here that we train a single MLP integrity monitor to use across all environments and conditions. This is achieved through a training dataset representative of both environments, collected in adverse conditions: we use simulation data for Office, and a section of the campus spatially separated from all test data for Campus. Across all our collected data, we include 13 out of the 23 total ‘difficult’ VPR instances put forward by [19]: pure vision matching; online processing; dense sampling/continuum of places; hard decisions on matching; a real-time system; storage limitations; dynamic environments with visual aliasing; sun glare and inconsistent illumination; in-sequence changing conditions; limited (or no) knowledge of conditions; changing velocity; reference set loops; and vehicle stoppage(s). We provide examples in Figure 5.
@@
The data capture process proceeded as follows: 1) a preliminary traverse of the environment was performed that maps the route to improve the accuracy of our analysis; then, 2) a reference traverse was performed which samples images and stores them with a position; finally, 3) a query traverse was performed which samples images and performs VPR. For Office and Campus, steps 2) and 3) were repeated to form the Normal and Adverse conditions. To limit lateral path error, traverses retraced the route from 1) using the ground truth.
@@ IV-F Multi-Layer Perceptron Hyperparameters
Throughout the described experiments, we train a single MLP integrity monitor for each VPR technique using a shared structure. We set the number of layers to $4$, the number of neurons per layer to $128$, and the output layer to a single neuron. For training purposes, we use a batch size of $8$, learning rate of $0.00001$, and a dropout rate of $10\%$. Due to differences in the training label balance when using different VPR techniques (out-of-tolerance: AP-GeM $=43.6\%$, NetVLAD $=47.2\%$, SALAD $=13.3\%$), we set $\alpha=6$, $\alpha=3$, $\alpha=35$, for these VPR techniques respectively.
@@ IV-G Comparisons
We compare the performance of our proposed system against baseline VPR with no verification, as well as the Support Vector Machine (SVM) integrity monitor from our previous work [44]. Additionally, as a diagnostic indicator, we compare against naive thresholds which reject VPR matches above a specified match distance, to attain a performance characteristic derived from our proposed system:
  • $N_{P}$: Equal precision to our proposed system.
  • $N_{R}$: Equal recall to our proposed system.
These thresholds are calculated during training for each VPR technique, and are consistent across all environments and conditions. They correspond to: $91.6\%$  precision, $21.3\%$  recall for AP-GeM; $88.2\%$  precision, $29.3\%$  recall for NetVLAD; and $98.6\%$  precision, $63.6\%$  recall for SALAD.

### v2

# Improving Visual Place Recognition Based Robot Navigation By Verifying Localization Estimates

@@ Abstract
Visual Place Recognition (VPR) systems often have imperfect performance, affecting the ‘integrity’ of position estimates and subsequent robot navigation decisions. Previously, SVM classifiers have been used to monitor VPR integrity. This research introduces a novel Multi-Layer Perceptron (MLP) integrity monitor which demonstrates improved performance and generalizability, removing per-environment training and reducing manual tuning requirements. We test our proposed system in extensive real-world experiments, presenting two real-time integrity-based VPR verification methods: a single-query rejection method for robot navigation to a goal zone (Experiment 1); and a history-of-queries method that takes a best, verified, match from its recent trajectory and uses an odometer to extrapolate a current position estimate (Experiment 2). Noteworthy results for Experiment 1 include a decrease in aggregate mean along-track goal error from $\approx\!9.8m$ to $\approx\!3.1m$ , and an increase in the aggregate rate of successful mission completion from $\approx\!41\%$ to $\approx\!55\%$. Experiment 2 showed a decrease in aggregate mean along-track localization error from $\approx\!2.0m$ to $\approx\!0.5m$, and an increase in the aggregate localization precision from $\approx\!97\%$ to $\approx\!99\%$. Overall, our results demonstrate the practical usefulness of a VPR integrity monitor in real-world robotics to improve VPR localization and consequent navigation performance.
Index Terms:
Localization; Acceptability and Trust; Vision-Based Navigation
@@ I Introduction
Figure 1: Overview of our system, demonstrating how our addition of a predictive verification system (a Multi-Layer Perceptron (MLP) network) results in safer navigation. See Figure 2 for more.
@@
    We present a novel approach using a Multi-Layer Perceptron network for VPR integrity monitoring, rather than the previously adopted Support Vector Machine (SVM), improving performance, generalizability across datasets, and outcomes in robot navigation paradigms.
@@
The letter proceeds as follows: in Section II we review the literature, in III we detail our approach, in IV we introduce our experimental methods, and in V we discuss the results.
@@ II-A Visual Place Recognition
Typically VPR involves extracting features from a current (query) image and comparing these with features from a set of reference images taken along a previously-traversed path to determine a best-match[3]. Numerous VPR techniques exist, with feature extraction and matching varying in complexity and performance [5, 6, 7, 8, 9, 10, 11], noting the recent techniques MixVPR [12], DinoV2 SALAD [13], and AnyLoc [14].
When deployed on real-world navigating robots, VPR is susceptible to potentially large localization errors due to varied environmental conditions, lighting, and presence of dynamic objects [15]. Although stand-alone systems have no match validation, some research has attempted to characterize the uncertainty. Uncertainty is often modelled using known distributions [16], however, other work shows VPR errors are often discontinuous and non-Gaussian [17]. Complementary to this, [18] had some success correlating VPR accuracy (within 25m) with network outputs by re-framing VPR as a classification problem. Sequence matching [19, 20, 21], Bayesian techniques [22, 23, 24] and particle filters [25, 26] may be used to help reduce error but adds cost in complexity and sensing. Some research has focused on predicting which VPR combinations will perform best in a given environment [27, 28], but these methods are computationally intensive.
@@ II-B Localization Integrity
Some integrity monitoring approaches have been proposed for stereo visual odometry [30, 31, 32, 33]; or for fusing visual odometry with sensors such as GNSS, LiDAR and radar [34, 35, 36, 37]. Other approaches include vision-based lane detectors [38], landmarks localization [39], or objects of interest localization within a scene [35]. However, none of these methods specifically address VPR. In our previous work [40], we proposed SVM-based integrity prediction methods for VPR such that each match’s accuracy is predicted to be within a tolerance; but this was investigated passively on datasets, not actively navigating robots. Regardless of method, integrating an integrity monitor is attractive as it can enable a robot to introspectively improve real-time operational decisions.
@@ III Approach
In this section we describe the proposed single-query and history-of-queries (HoQ) localization methods.
@@ III-A VPR Technique Requirements
The integrity monitor we build on from [40] only requires some form of distance vector input for predictions, making it compatible with many, if not all, VPR techniques. The distance vectors consist of mathematical distances between query image features and those from a reference set; we refer to these distances as ‘match distances’. In this work, we maintain compatibility with all, or at least the vast majority of, VPR techniques by similarly using information already available from the VPR process to verify VPR integrity. We demonstrate this by testing using three VPR techniques: AP-GeM [41], NetVLAD [5], and SALAD [13].
@@ III-B Integrity Prediction
We improve on the previously adopted SVM model from [40] by implementing a small multi-layer perceptron (MLP) network, with a simple fully-connected structure, to utilise the increased learning capability of neural networks whilst maintaining the light-weight nature of the predictor. We use 48 hand-crafted statistical feature calculations which, when applied to four vectors, yields a total of 192 unique features as the classification input. For brevity we include a simplified diagram of the process (Fig. 2); the detailed mathematical formulation and the software implementation is provided in the mentioned repository1. We use such a large number of statistical features to remove the necessity of knowing the best features for individual VPR techniques.
Figure 2: An overview of how inputs for the MLP integrity monitor are extracted. For a given query, we use four vectors formed from the output of the VPR process. The vectors pass through a statistical feature extractor and are then concatenated. These steps guarantee the size of the input is suitable for the MLP network, which accepts a 192-dimensional vector.
@@ III-C MLP Integrity Monitor Training
For a given query image with index $k$, the input vector for the MLP network becomes a 192-dimensional vector created through concatenating four groups of 48 statistical features. These groups are calculated from the $k^{\text{th}}$: distance vector, $\mathbf{D}$ [^7] Bold denotes arrays. Subscripts denote indices, which start from $1$., as per [40]; the query VPR feature vector, $\mathbf{Q}$; the best-match reference feature vector, $\mathbf{R}$; and the VPR feature difference vector, $\mathbf{V}=\mathbf{R}-\mathbf{Q}$. The query training label, $\mathbf{P}_{k}$, simply becomes a binary value indicating whether the top VPR match corresponds to a reference in-tolerance, $1$, or out-of-tolerance, 00. We formulate the problem as a regression task using a weighted mean-squared-error (MSE) loss, providing control over the rate queries are predicted as out-of-tolerance. The loss function, $L(\mathbf{P},\mathbf{\hat{P}},\alpha)$, is defined as:
@@
where $\mathbf{\hat{P}}_{k}$ is the integrity prediction from the MLP network, $\alpha$ is a scalar weight, and $N$ is the dimension of $\mathbf{P}$ and $\mathbf{\hat{P}}$. We set $\alpha>1$ to produce a cautious integrity monitor that reduces the number of out-of-tolerance VPR matches that are incorrectly predicted as in-tolerance (false positives), given these pose a greater risk to robot navigation tasks. Effectively, $\alpha$ controls the precision and recall of classification performance for the integrity prediction task.
III-D Localizing with VPR and Integrity: Single-Query Method
In the single-query scenario where each query is treated independently, the VPR system performs query-to-reference matching and provides a position estimate using the retrieved 2D pose ($x$,$y$,$\theta$) from where the best-match reference image was collected. The integrity monitor (Section III-B) then verifies the unprocessed output by only accepting position estimates that are predicted to be in-tolerance of the true position; all predicted out-of-tolerance estimates are rejected.
Figure 3: HoQ method: we rank a recent history of VPR matches by their match distance (1), reduce to verified matches (2), then extrapolate from the best, verified match to estimated position using the odometer difference (3).
III-E Utilizing Historical Data: History-of-Queries Method
We can improve the single-query localization decisions (Section III-D) by utilizing an odometer to extrapolate a position estimate based on a recent history of predicted in-tolerance VPR matches, which we refer to as the history-of-queries (HoQ) method. The use of historical information is a well-known component of many navigation systems: our key contribution here is the added precursor step where all entries in the history are verified (such that only predicted in-tolerance matches are retained), distinguishing our method from those which only perform statistical or threshold filtering. If the system finds at least one verified historical match, then it proceeds to localize. We diagram this in Figure 3.
This process can be described by the following. Consider some unknown current vehicle pose $\chi$ and known odometer scalars $\omega$. Within the previous $d$ meters, we sample VPR match indices $\mathbf{M}$ and odometer scalars $\mathbf{W}$, collected together at irregular spatiotemporal intervals. For each match index $\mathbf{M}_{i}$, we find a corresponding: match distance $\mathbf{D}_{i}$, integrity prediction $\mathbf{\hat{P}}_{i}$, and pose estimate $\mathbf{X}_{\mathbf{M}_{i}}\in\mathbf{X}$, where $\mathbf{X}$ is the ordered set of 2D poses with elements indexed to each reference image (dimension $n$). We assume $\chi\equiv\mathbf{X}_{q}$, where $q$ is the index of the closest (Euclidean) pose in $\mathbf{X}$ to $\chi$. Using $\mathbf{M},\mathbf{D},\mathbf{W},\mathbf{\hat{P}}$, and $\mathbf{X}_{\mathbf{M}}$ as the recent history contents, per Figure 3, we now work to determine $q$.
@@ III-E Utilizing Historical Data: History-of-Queries Method
We then compute the running sum of the elements of $\mathbf{T}$, denoted as $\mathbf{S}$, from which we find $q$ and $\chi$ per Fig. 3, Step 3:
@@ IV Experimental Procedure
In this section, we introduce our two experiment designs and items such as hardware, datasets, parameters, and metrics.
@@ IV-A Experiment 1 Design
In the first experiment scenario, a robot moves along a path whilst attempting to autonomously determine if it has reached an end-goal zone. We implement the single-query method introduced in Section III-D for localization. The mission is successfully completed if the robot correctly identifies it has arrived at the end-goal within an assessment tolerance of $\pm$0.5m, otherwise the mission is failed, as shown in Fig. 4.
@@
For fair assessment, given the underlying baseline system performance changes along the route due to environmental and dynamic variations, we randomly select 50 start locations over the reference path, with end-goal locations allocated at 5m, 10m, 25m and 50m from each starting position. As navigation decisions require a tighter tolerance, the robot will flag arrival if its localization estimate is anywhere beyond the point 10cm before the end-goal center. This provides a margin of error in achieving the looser, actual, target tolerance. We emphasize that VPR localizes from a reference set of the whole environment, not the subset being traversed.
@@ IV-B Experiment 2 Design
In the second experiment scenario, a robot repeats a pre-mapped traverse whilst localizing using the HoQ method from Section III-E. In this scenario, the goal is to maximize precision. We select a history length of 1.5m. For fair analysis the robot must initialize the history by traversing this distance.
@@ IV-D Data Collection: Environments and Conditions
We conducted both experiments in two different environments: Office, containing indoor images from the QUT Centre for Robotics (QCR); and Campus, containing outdoor images from the QUT Gardens Point campus. For both environments we recorded two sets of conditions: Normal, where there is a typical presence of adversity given the environment (some minor lighting variation and dynamic or static objects); and Adverse, where there is a heightened presence of adversity, such as increased lighting variation (morning to evening), and more dynamic or static objects. Data capture periods were selected to control these introduced adversities. We emphasize here that we train a single MLP integrity monitor to use across all environments and conditions. This is achieved through a representative training dataset of each environment, formed from two query and reference pairs, captured under Adverse conditions. For these we use simulation data for Office, and a section of the campus spatially separated from all test data for Campus. Across all our collected data, we include 13 out of the 23 total ‘difficult’ VPR instances put forward by [15]; we provide examples in Figure 5.
@@
For simplicity in analysing performance in this novel setting, lateral path error was limited in traverses by retracing the reference route using ground truth. In real scenarios, VPR must compensate for lateral shift, which will likely reduce the precision of integrity monitors. This could be somewhat mitigated by controlling training data to include representative samples with viewpoint shifts, however, we believe the additional information and statistical features in the proposed MLP approach should improve robustness over previous SVM approaches.
@@ IV-F Multi-Layer Perceptron Hyperparameters
Throughout the described experiments, we train a single MLP integrity monitor for each VPR technique using a shared structure. We set the number of layers to $4$, the number of neurons per layer to $128$, the output layer to a single neuron, with ReLU activations between layers and finish with a Sigmoid activation. For training purposes, we use a batch size of $8$, learning rate of $0.00001$, and a dropout rate of $10\%$. Due to differences in the training label balance when using different VPR techniques (out-of-tolerance: AP-GeM $=43.6\%$, NetVLAD $=47.2\%$, SALAD $=13.3\%$), we set $\alpha=6$, $\alpha=3$, $\alpha=35$, for these VPR techniques respectively.
@@ IV-G Comparisons
We compare the performance of our proposed system against baseline VPR with no verification, as well as the SVM integrity monitor from our previous work [40]. Additionally, we compare against naive thresholds which reject VPR matches above a specified match distance, as diagnostic indicators for the performance characteristics of our proposed system: $N_{P}$, with equal precision to our proposed system; and $N_{R}$, with equal recall to our proposed system. These thresholds are calculated during training for each VPR technique, and are consistent across all environments and conditions. They correspond to: $91.6\%$  precision, $21.3\%$  recall for AP-GeM; $88.2\%$  precision, $29.3\%$  recall for NetVLAD; and $98.6\%$  precision, $63.6\%$  recall for SALAD.